A questionnaire asks shareholders of a company to state whether they consider the chairman’s salary to be too high, about right, or too low. Excluding shareholders who have no opinion, the probabilities of answers from randomly selected shareholders are as follows:</span><span lang="EN-US"><o:p></o:p></span>

Too High       0.80

About Right   0.25

Too Low         0.05

           What are the probabilities that if three shareholders are selected at random,

a.      They will all answer, ‘too high’?

b.     Exactly two will answer ‘too low’?

c.      Exactly two will give the same answer?


1
Expert's answer
2022-03-17T07:05:42-0400

"a:P=0.8\cdot 0.8\cdot 0.8=0.512\\b:P=C_{3}^{2}\cdot 0.05^2\cdot \left( 1-0.05 \right) =0.007125\\c:P=C_{3}^{2}\cdot \left( 0.8^2\left( 1-0.8 \right) +0.25^2\left( 1-0.25 \right) +0.05^2\left( 1-0.05 \right) \right) =0.53175"